4x – 3y = –10
7x + 3y = –12

Answer by Boreal(15235)

4x – 3y = –10
7x + 3y = –12 When you see that adding them will make the y disappear, it becomes a lot easier.
11x=-22
x=-2
substitute in the first (or second)
4(-2)-3y=-10
-8-3y=-10
-3y=-2
y=2/3
Check in second equation
7(-2)+3(2/3)=-14+2=-12, checks
(-2,2/3) ANSWER